OpenStudy (anonymous):

what is the integral along c of following Re z dz where c is the shortest path from 1+i to 3+2i.

OpenStudy (anonymous):

I suggest drawing a picture

OpenStudy (anonymous):

i didn't get what you said

OpenStudy (anonymous):

solve it using cauchy integral formula or theorem etc.

OpenStudy (mr.math):

The first thing to do here is to find the parametrization of the given curve, which is given by: \(z(t)=1+i+t(3+2i-1-i)=1+i+t(2+i)\).

OpenStudy (mr.math):

For the parametrization above: \(0\le t\le 1\)

OpenStudy (mr.math):

You probably know the formula: \[\int\limits_{\zeta}f(z)dz=\int\limits_{a}^{b}f(z(t))z'(t)dt\] where \(z(t)\) is the parametrization of \(\zeta\).

OpenStudy (mr.math):

Now all we need is a direct substitution with \(z(t)=1+i+t(2+i)\), \(z'(t)=2+i\), \(a=0\) and \(b=1\).

OpenStudy (anonymous):

thanks alot plz answer my second question